### v1

# Homeomorphism theorem for sums of translates on the real axis

Abstract.
In this paper, we study sums of translates on the real axis. These functions generalize logarithms of weighted algebraic polynomials. Namely, we are dealing with the following functions
$F(\mathbf{y},t):=J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j}),\quad\mathbf{y}:=(y_{1},\ldots,y_{n}),\ y_{1}\leqslant\ldots\leqslant y_{n},$
where the field function $J$ is a function defined on $\mathbb{R}$, which is "admissible" for the kernels $K_{1},\ldots,K_{n}$ concave on $(-\infty,0)$ and on $(0,\infty)$ and having a singularity at $0.$ We consider "local maxima"
$\begin{aligned} m_{0}(\mathbf{y})&:=\sup\limits_{t\in(-\infty,y_{1}]}F(\mathbf{y},t),\quad m_{n}(\mathbf{y}):=\sup\limits_{t\in[y_{n},\infty)}F(\mathbf{y},t),\\ m_{j}(\mathbf{y})&:=\sup\limits_{t\in[y_{j},y_{j+1}]}F(\mathbf{y},t),\quad j=1,\ldots,n-1,\end{aligned}$
and the difference function
$D(\mathbf{y}):=(m_{1}(\mathbf{y})-m_{0}(\mathbf{y}),m_{2}(\mathbf{y})-m_{1}(\mathbf{y}),\ldots,m_{n}(\mathbf{y})-m_{n-1}(\mathbf{y})).$
We prove that, under certain assumptions on monotonicity of the kernels, $D$ is a homeomorphism between its domain and $\mathbb{R}^{n}.$
Key words and phrases:
Sums of translates function, locally bi-Lipschitz homeomorphism, homeomorphism theorem, abstract $\log$-concave interpolation, moving node Hermite–Fejér interpolation
2010 Mathematics Subject Classification:
41A50, 41A52, 42A15, 26A51
The work was performed as part of research conducted in the Ural Mathematical Center with the financial support of the Ministry of Science and Higher Education of the Russian Federation (Agreement number 075-02-2025-1549).
1. Introduction
In this paper, we study sums of translates on the real axis. These functions generalize logarithms of weighted algebraic polynomials. Namely, we are dealing with the following functions
$F(\mathbf{y},t):=J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j}),\quad\mathbf{y}:=(y_{1},\ldots,y_{n}),\ y_{1}\leqslant\ldots\leqslant y_{n},$
where the field function $J$ is a function defined on $\mathbb{R}$, which is "admissible" for the kernels $K_{1},\ldots,K_{n}$ concave on $(-\infty,0)$ and on $(0,\infty)$ and having a singularity at $0.$
The sums of translates and the minimax problem for such functions were first considered by P. C. Fenton in 2000 [1]. He considered one kernel with assumptions of monotonicity, smoothness, singularity of its derivative at 0, and a concave field $J$ continuous at the ends of the segment. Fenton’s original goal was to prove P. D. Barry’s conjecture from 1962 on the growth of entire functions, which he succeeded in 1981 [2]. And although A. A. Goldberg proved this conjecture a little earlier [3], Fenton obtained other interesting results in the theory of entire functions using his approach [4], [5].
In our previous paper [6], we studied the minimax problem for the sums of translates on the real axis. The main method was a reduction to the minimax theorem proved by B. Farkas, B. Nagy and Sz. Gy. Révész for a segment (see [7], [8]). The uniqueness of the minimax point followed immediately from this reduction. To prove the uniqueness of the minimax point on the segment, the authors used the so-called homeomorphism theorem [9, Th. 7.1], which was also proved by them. Now, we prove a similar homeomorphism theorem for the real axis. As for the segment, this result provides the uniqueness of the minimax point on the real axis, independently of the specific reduction technique.
We consider "local maxima"
(1) $m_{0}(\mathbf{y})$ $=\sup\limits_{t\in(-\infty,y_{1}]}F(\mathbf{y},t),\quad m_{n}(\mathbf{y})=\sup\limits_{t\in[y_{n},\infty)}F(\mathbf{y},t),$
$m_{j}(\mathbf{y})$ $=\sup\limits_{t\in[y_{j},y_{j+1}]}F(\mathbf{y},t),\quad j=1,\ldots,n-1,$
and the difference function
$D(\mathbf{y}):=(m_{1}(\mathbf{y})-m_{0}(\mathbf{y}),m_{2}(\mathbf{y})-m_{1}(\mathbf{y}),\ldots,m_{n}(\mathbf{y})-m_{n-1}(\mathbf{y})).$
We prove that, under certain assumptions on monotonicity of the kernels, $D$ is a homeomorphism between its domain and $\mathbb{R}^{n}.$
Results of this kind are inspired by the problem of optimizing the Lagrange interpolation of a continuous function. Let us give an overview of the results known to us.
Let $f$ be a function continuous on $[0,1]$ and $y_{0}=0<y_{1}<\ldots<y_{n}<1=y_{n+1}$ be interpolation nodes. Denote by $\pi_{n+1}$ the space of polynomials of degree at most $n+1$ and by $P_{\mathbf{y}}:C_{[0,1]}\to\pi_{n+1}$ the Lagrange interpolation operator
$P_{\mathbf{y}}f(t):=\sum\limits_{j=0}^{n+1}f(y_{j})\ell_{j}(t),\quad\ell_{j}(t):=\prod\limits_{i\neq j}\dfrac{t-y_{j}}{y_{i}-y_{j}}.$
It is easy to show that
$\|P_{\mathbf{y}}f-f\|_{C_{[0,1]}}\leqslant\operatorname{dist}(f,\pi_{n+1})(1+\|P_{\mathbf{y}}\|),$
where $\|P_{\mathbf{y}}\|$ is the operator norm. Therefore, it is natural to minimize $\|P_{\mathbf{y}}\|$ by $\mathbf{y}$ to optimize the interpolation. It is known that $\|P_{\mathbf{y}}\|=\|\Lambda_{\mathbf{y}}\|_{C_{[0,1]}},$ where $\Lambda_{\mathbf{y}}(t):=\sum\limits_{j=0}^{n+1}|\ell_{j}(t)|$ is the Lebesgue function. Thus, this optimization problem is exactly a minimax problem for the functions $\Lambda_{\mathbf{y}}.$ Denote $\lambda_{j}(\mathbf{y}):=\max\limits_{t\in[y_{j},y_{j+1}]}\Lambda_{\mathbf{y}}(t),\ j=0,\ldots,n.$
In 1931, S. N. Bernstein [10] conjectured that the minimum of $\|\Lambda_{\mathbf{y}}\|_{C_{[0,1]}}$ is attained when $\mathbf{y}$ is an equioscillation point, i. e.,
$\lambda_{0}(\mathbf{y})=\ldots=\lambda_{n}(\mathbf{y}).$
In 1977, T. A. Kilgore [11] proved Bernstein’s conjecture. Moreover, he showed the uniqueness of the equioscillation point. More precisely, Kilgore’s note [12] describing the proof of the statement “the minimax point is an equioscillation point” was first published, and a few months later, the complete proof was presented in [11]. Subsequently, C. R. de Boor and A. Pinkus [13] also proved Bernstein’s conjecture. In fact, they obtained the following general result.
Theorem A.
Let $S^{[0,1]}:=\{\mathbf{y}=(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}:\ 0<y_{1}<\ldots<y_{n}<1\}.$ Then the difference function
$D_{\lambda}:S^{[0,1]}\to\mathbb{R}^{n},\quad\mathbf{y}\mapsto(\lambda_{1}(\mathbf{y})-\lambda_{0}(\mathbf{y}),\ldots,\lambda_{n}(\mathbf{y})-\lambda_{n-1}(\mathbf{y}))$
is a homeomorphism between $S^{[0,1]}$ and $\mathbb{R}^{n}.$
It immediately implies that there is exactly one equioscillation point. De Boor and Pinkus then refer to Kilgore’s earlier note with the proof that the minimax point is an equioscillation point, and together all this proves Bernstein’s conjecture.
In 1996, Y. G. Shi [14] considered more general functions $\varphi_{j}(\cdot),\ j=0,\ldots,n,$ instead of $\lambda_{j}(\cdot).$ Shi supposed that all the functions $\varphi_{j}$ are continuously differentiable on $S^{[0,1]}$ and satisfy conditions
$\min\limits_{0\leqslant j\leqslant n}\lim\limits_{y_{j+1}-y_{j}\to 0}\ \max\limits_{0\leqslant i\leqslant n-1}|\varphi_{i+1}(\mathbf{y})-\varphi_{i}(\mathbf{y})|=\infty$
and
$\Phi_{k}(\mathbf{y}):=\det\left(\dfrac{\partial\varphi_{i}(\mathbf{y})}{\partial y_{j}}\right)^{n\ n}_{j=1,\ i=0,\ i\neq k}\neq 0,\quad\mathbf{y}\in S^{[0,1]},\ k=0,\ldots,n.$
Shi dealt with the following minimax problem: find a vector $\mathbf{y}=(y_{1},\ldots,y_{n}),\ 0<y_{1}<\ldots<y_{n}<1,$ that minimizes $\max\limits_{j=0,\ldots,n}\varphi_{j}(\cdot).$ Under these assumptions, Shi proved that there exists a unique extremal point and it has the equioscillation property. Moreover, Shi obtained the homeomorphism theorem for the difference function
$D_{\varphi}:S^{[0,1]}\to\mathbb{R}^{n},\quad\mathbf{y}\mapsto(\varphi_{1}(\mathbf{y})-\varphi_{0}(\mathbf{y}),\ldots,\varphi_{n}(\mathbf{y})-\varphi_{n-1}(\mathbf{y})).$
In particular, this theorem implies the uniqueness of the minimax point.
In 2018, B. Farkas, B. Nagy and Sz. Gy. Révész presented a solution of the minimax problem and a homeomorphism result for sums of translates $F(\mathbf{y},t):=K_{0}(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j})$ on a torus [15]. Here $K_{0},\ldots,K_{n}:\ \mathbb{R}\to[-\infty,0)$ are $2\pi$-periodic functions, strictly concave on $(0,2\pi)$. Assuming that for each $j=0,\ldots,n$ the function $K_{j}$ belongs to $C^{2}(0,2\pi)$ with $K^{\prime\prime}_{j}<0$ and $K(0)=K(2\pi)=-\infty,$ they proved that the difference function of local maxima is a homeomorphism between its domain and $\mathbb{R}^{n}$. On the one hand, the sums of translates approach is more specific than Shi’s. On the other hand, the authors provided an example [15, Ex. 5.13] demonstrating that Shi’s result is not applicable in their settings.
In 2021, Farkas, Nagy and Révész proved a homeomorphism theorem for sums of translates on the segment [9, Th. 7.1]. This result is the basis of our proof, since to solve the problem on the real axis we reduce it to the case of the segment. For a precise formulation of the problems, we need a little more preparation.
Definition 1.1.
Let $0<p\leqslant\infty.$ A function $K\colon(-p,0)\cup(0,p)\to\mathbb{R}$ is called a kernel function if $K$ is concave on (p,0)(-p,0) and on $(0,p)$ and $\lim\limits_{t\downarrow 0}K(t)=\lim\limits_{t\uparrow 0}K(t)$, which are either real or equal to -$\infty$.
We extend $K$ by defining
$K(0):=\lim\limits_{t\to 0}K(t),\quad K(-p):=\lim\limits_{t\downarrow-p}K(t),\quad K(p):=\lim\limits_{t\uparrow p}K(t).$
If $K(0)=$-$\infty$, the kernel function $K$ is called singular.
When $K$ is defined on $(-\infty,0)\cup(0,\infty),$ the following condition, called generalized monotonicity, is important to us:
(GMGM) $\lim\limits_{t\to-\infty}K^{\prime}(t)\leqslant\lim\limits_{t\to\infty}K^{\prime}(t).$
By the concavity of $K$, the set where $K^{\prime}$ is defined has full measure. We consider the limits of $K^{\prime}$ on this set.
Remark 1.1.
[…]
If $K$ is (strictly) decreasing on (p,0)(-p,0) and (strictly) increasing on $(0,p)$, then we call $K$ (strictly) monotone. Obviously, if $K$ is defined on $(-\infty,0)\cup(0,\infty)$ and monotone, then it satisfies ($GM$).
Remark 1.2.
There is a direct connection between the notions of monotonicity and generalized monotonicity ($GM$). Specifically, condition ($GM$) is equivalent to the existence of a number $c$ such that the kernel K(t)ctK(t)-ct is monotone.
Indeed, suppose that (GMGM) holds. Then, defining $c:=\lim\limits_{t\to\infty}K^{\prime}(t)$, we consider the function $\widetilde{K}(t):=K(t)-ct$. By (GMGM), its derivative satisfies
$\lim\limits_{t\to-\infty}\widetilde{K^{\prime}}(t)=\lim\limits_{t\to-\infty}K^{\prime}(t)-c\leqslant 0\quad\text{and}\quad\lim\limits_{t\to\infty}\widetilde{K^{\prime}}(t)=\lim\limits_{t\to\infty}K^{\prime}(t)-c=0.$
Since $K^{\prime}$ is non-decreasing at all points of its domain, this implies that $\widetilde{K^{\prime}}$ is monotone.
Conversely, if there exists a number $c$ such that the kernel K(t)ctK(t)-ct is monotone, then its derivative satisfies
$\lim\limits_{t\to-\infty}K^{\prime}(t)-c\leqslant 0\quad\text{and}\quad\lim\limits_{t\to\infty}K^{\prime}(t)-c\geqslant 0.$
Hence, obviously, (GMGM) holds.
Definition 1.2.
Let $A$ be a segment, a semiaxis or $\mathbb{R}$. We call a function $J\colon A\to\underline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty\}$ an external $n$-field function or simply a field on $A$ if $J$ is bounded above on $A$ and it assumes finite values at more than $n$ different points of $A,$ where in the case of a segment we count boundary points with weights $1/2.$
In the case of a segment, it is necessary that there are at least $n$ interior points and some additional one anywhere in the segment, where the field is finite. Therefore, we impose precisely such conditions on the weights of the points to ensure consistency with the case of a segment.
Consider a segment $[a,b]$. In what follows, we denote by $\overline{S^{[a,b]}}$ the closed simplex
$\overline{S^{[a,b]}}:=\{(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}:\ a\leqslant y_{1}\leqslant\ldots\leqslant y_{n}\leqslant b\}.$
Consider the sums of translates
$F(\mathbf{y},t):=J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j}),\quad\mathbf{y}\in\overline{S^{[a,b]}},\ t\in[a,b].$
Denote
$m^{[a,b]}_{0}(\mathbf{y})$ $:=\sup\limits_{t\in[a,y_{1}]}F(\mathbf{y},t),\quad m^{[a,b]}_{n}(\mathbf{y}):=\sup\limits_{t\in[y_{n},b]}F(\mathbf{y},t),$
$m^{[a,b]}_{j}(\mathbf{y})$ $:=\sup\limits_{t\in[y_{j},y_{j+1}]}F(\mathbf{y},t),\quad j=1,\ldots,n-1.$
Let us introduce the regularity set
$R^{[a,b]}:=\{\mathbf{y}\in\overline{S^{[a,b]}}:\ m^{[a,b]}_{j}(\mathbf{y})\neq-\infty\text{ for }j=0,\ldots,n\}$
and the difference function $D^{[a,b]}:R^{[a,b]}\to\mathbb{R}^{n}$
$D^{[a,b]}(\mathbf{y}):=(m^{[a,b]}_{1}(\mathbf{y})-m^{[a,b]}_{0}(\mathbf{y}),m^{[a,b]}_{2}(\mathbf{y})-m^{[a,b]}_{1}(\mathbf{y}),\ldots,m^{[a,b]}_{n}(\mathbf{y})-m^{[a,b]}_{n-1}(\mathbf{y})).$
If b=ab=-a, then we write $\overline{S^{b}},\ m^{b}_{j}(\mathbf{y}),\ R^{b}.$
As we mentioned above, the following homeomorphism theorem was proven by Farkas, Nagy and Révész [9, Th. 7.1].
Theorem B.
Let $a<b,$ the kernel functions $K^{[a,b]}_{1},\ldots,K^{[a,b]}_{n}:(a-b,0)\cup(0,b-a)\to\mathbb{R}$ be singular, strictly concave, and $J^{[a,b]}:[a,b]\to\underline{\mathbb{R}}$ be an $n$-field function. Assume that
(2) $\left(K^{[a,b]}_{j}(t)-K^{[a,b]}_{j}(t-(b-a))\right)^{\prime}\geqslant 0\quad\text{for almost all}\quad t\in(0,b-a),\quad j=1,\ldots,n,$
and
(3) $J^{[a,b]}(a)=\lim\limits_{t\downarrow a}J^{[a,b]}(t)=-\infty\quad\text{or}\quad J^{[a,b]}(b)=\lim\limits_{t\uparrow b}J^{[a,b]}(t)=-\infty.$
Then the difference function
$D^{[a,b]}:R^{[a,b]}\to\mathbb{R}^{n},\quad\mathbf{y}\mapsto(m^{[a,b]}_{1}(\mathbf{y})-m^{[a,b]}_{0}(\mathbf{y}),m^{[a,b]}_{2}(\mathbf{y})-m^{[a,b]}_{1}(\mathbf{y}),\ldots,m^{[a,b]}_{n}(\mathbf{y})-m^{[a,b]}_{n-1}(\mathbf{y}))$
is a homeomorphism between $R^{[a,b]}$ and $\mathbb{R}^{n}$. Moreover, $D^{[a,b]}$ is locally bi-Lipschitz.
Remark 1.3.
The authors proved the homeomorphism theorem for a=0,b=1.a=0,\ b=1. We formulate this theorem for an arbitrary segment for convenience for our later application. The theorem on $[0,1]$ can be trivially extended to $[a,b]$ by applying the linear transformation
$\chi(t):=\dfrac{t-a}{b-a}$
that maps the segment $[a,b]$ onto $[0,1]$. Let us show this.
Let $K^{[a,b]}_{j},\ j=1,\ldots,n,$ be kernel functions defined on $(a-b,0)\cup(0,b-a)$ and $J^{[a,b]}$ be a field function on $[a,b]$. Assume that these kernels and field satisfy the conditions of Theorem B. Let us show that the homeomorphism theorem proven on $[0,1]$ implies Theorem B. Let
$\mathscr{K}^{[0,1]}_{j}(x)$ $:=K^{[a,b]}_{j}((b-a)x),\quad x\in[-1,1],\ j=1,\ldots,n,$
$\mathscr{J}^{[0,1]}(x)$ $:=J^{[a,b]}(a+(b-a)x),\quad x\in[0,1].$
We have for $t\in[a,b]$ that
(4) $K^{[a,b]}_{j}(t-y_{j})$ $=\mathscr{K}^{[0,1]}_{j}(\chi(t)-\chi(y_{j})),\quad j=1,\ldots,n,$
$J^{[a,b]}(t)$ $=\mathscr{J}^{[0,1]}(\chi(t)).$
Hence, taking into account the conditions on $K^{[a,b]}_{j}$ and $J^{[a,b]}$, we have that the kernels $\mathscr{K}^{[0,1]}_{j}$ are strictly concave and singular and $\mathscr{J}^{[0,1]}$ is an $n$-field function satisfying
$\mathscr{J}^{[0,1]}(0)=\lim\limits_{x\downarrow 0}\mathscr{J}^{[0,1]}(x)=-\infty\quad\text{or}\quad\mathscr{J}^{[0,1]}(1)=\lim\limits_{x\uparrow 1}\mathscr{J}^{[0,1]}(x)=-\infty.$
Let us show that
(5) $\left(\mathscr{K}^{[0,1]}_{j}(x)-\mathscr{K}^{[0,1]}_{j}(x-1)\right)^{\prime}_{x}\geqslant 0\quad\text{for almost all}\quad x\in(0,1).$
We have
$\left(K^{[a,b]}_{j}(t)-K^{[a,b]}_{j}(t-(b-a))\right)^{\prime}_{t}=\left(\mathscr{K}^{[0,1]}_{j}(\chi(t)-\chi(0))-\mathscr{K}^{[0,1]}_{j}(\chi(t)-\chi(b-a))\right)^{\prime}_{t}$
$=\left(\mathscr{K}^{[0,1]}_{j}(t/(b-a))-\mathscr{K}^{[0,1]}_{j}(t/(b-a)-1)\right)^{\prime}_{t}.$
For $t\in(0,b-a)$, substituting $x:=t/(b-a)\in(0,1)$ and using (2), we obtain (5).
Therefore, the difference function $\mathscr{D}^{[0,1]}$ is a homeomorphism between $\mathscr{R}^{[0,1]}$ and $\mathbb{R}^{n}.$ Denote $\boldsymbol{\chi}(\mathbf{y}):=(\chi(y_{1}),\ldots,\chi(y_{n}))$. Using (4), it is easy to see that
$\boldsymbol{\chi}(R^{[a,b]})=\mathscr{R}^{[0,1]},\quad D^{[a,b]}(\mathbf{y})\equiv\mathscr{D}^{[0,1]}(\boldsymbol{\chi}(\mathbf{y})),\quad\mathbf{y}\in R^{[a,b]}.$
So, we obtain that $D^{[a,b]}$ is a homeomorphism, too.
In addition to [9, Th. 7.1], the authors obtained homeomorphism theorems with other conditions on the kernels and field. Conditions (3) may be replaced by so-called cusp conditions at the ends of the segment [9, Th. 7.5]. Moreover, if the derivatives of the kernel differences in (2) are bounded below by some $c>0,$ then the field can be arbitrary [9, Th. 2.1].
The minimax problem for the sums of translates on the segment was also deeply studied by Farkas, Nagy and Révész in [7], [8]. In particular, they proved that if the kernel $K^{[a,b]}$ is monotone, singular and strictly concave, then there exists a minimax point characterized by the equioscillation property. To prove the uniqueness of the equioscillation point, Farkas, Nagy and Révész apply Theorem B, which immediately implies this. Their research inspired the author to obtain similar results for the minimax problem on the real axis [6] and to write this paper.
Our goal is to prove an analogue of Theorem B for sums of translates on $\mathbb{R}$. Our method relies on reducing the problem to Theorem B, using an approach developed primarily in our previous paper [6]. Let us introduce the main definitions and state our result.
Let $K_{j}:(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be kernels and $J:\mathbb{R}\to\underline{\mathbb{R}}$ be an $n$-field. We consider the “infinite closed simplex”
$\overline{S^{\infty}}:=\{(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}\colon-\infty<y_{1}\leqslant y_{2}\leqslant\ldots\leqslant y_{n}<\infty\}.$
Similarly to the finite interval case, we can consider the regularity set
$R:=\{\mathbf{y}\in\overline{S^{\infty}}:\ m_{j}(\mathbf{y})\neq-\infty\text{ for }j=0,\ldots,n\}$
and also the corresponding difference function
(6) $D:R\to\mathbb{R}^{n},\ \mathbf{y}\mapsto(m_{1}(\mathbf{y})-m_{0}(\mathbf{y}),m_{2}(\mathbf{y})-m_{1}(\mathbf{y}),\ldots,m_{n}(\mathbf{y})-m_{n-1}(\mathbf{y})).$
Definition 1.3.
Let $K_{j}:(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be kernels. An $n$-field function $J$ defined on $\mathbb{R}$ is said to be admissible (for $K_{1},\ldots,K_{n}$) if
$\lim_{|t|\to\infty}\left(J(t)+\sum\limits_{j=1}^{n}K_{j}(t)\right)=-\infty.$
This condition is a version of the condition for admissible weights in weighted potential theory, see [16, p. 26]
We prove the following result.
Theorem 1.1.
Suppose that the singular, strictly concave kernel functions $K_{j}:(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ satisfy ($GM$) and $J:\mathbb{R}\to\underline{\mathbb{R}}$ is an admissible $n$-field function for $K_{1},\ldots,K_{n}$. Then the difference function defined in (6) is a homeomorphism between $R$ and $\mathbb{R}^{n}$. Moreover, $D$ is locally bi-Lipschitz.
2. Simple lemmas about kernels and fields
In this section, we need the following extension of the fundamental theorem of calculus for concave functions. It is known [17, p. 9] that if $g$ is a concave function on an interval $I$, then for any $a,b]\leavevmode\nobreak\ \subset\leavevmode\nobreak\ I$
(7) $g(b)-g(a)=\int\limits_{[a,b]}g^{\prime}(t)dt.$
Lemma 2.1.
Let $K:(-\infty,0)\cup(0,\infty)\to\mathbb{R}$ be a kernel function.
  1. (1)
    If $0<t_{1}<t_{2}<t_{2}+h$ or $t_{1}<t_{2}<t_{2}+h<0$, then
    (8) $K(t_{2}+h)-K(t_{1}+h)\leqslant K(t_{2})-K(t_{1}).$
    Moreover, $K^{\prime}$ is non-increasing at all points of its domain.
  2. (2)
    If $K$ satisfies (GMGM) and $t_{1}<t_{1}+h<0<t_{2}$, then
    (9) $K(t_{2})-K(t_{1})\leqslant K(t_{2}+h)-K(t_{1}+h).$
    In particular, for almost all $t_{1},t_{2}$ such that $t_{1}<0<t_{2}$ we have
    (10) $K^{\prime}(t_{1})\leqslant K^{\prime}(t_{2}).$
Proof.
  1. (1)
    Inequality (8) follows from definition of concavity. Its proof can be found e. g. in [18, Lemma 10].
    To prove the statement about $K^{\prime}$, it is sufficient to assume that $K^{\prime}$ is defined at $t_{2}$ and $t_{2}+h,$ divide (8) by $t_{2}-t_{1}$ and pass to the limit.
  2. (2)
    Sufficiently using (7), point 1 and (GMGM), we get
    $(K(t_{2}+h)-K(t_{2}))-(K(t_{1}+h)-K(t_{1}))=$
    $\int\limits_{[t_{2},t_{2}+h]}K^{\prime}(t)-\int\limits_{[t_{1},t_{1}+h]}K^{\prime}(t)\geqslant h\left(\lim\limits_{t\to\infty}K^{\prime}(t)-\lim\limits_{t\to-\infty}K^{\prime}(t)\right)\geqslant 0.$
    By Remark 1.1, the limits in (GMGM) are finite, so their difference is well-defined. Thus we have obtained (9).
    To get (10), one can group the terms in (9) with $t_{1}$ and $t_{2}$ together on the left respectively on the right hand side, then divide by $h$ and pass to the limit.
Let us show that admissibility of a field is equivalent to the following more general property.
Lemma 2.2.
Let $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be kernels. If a field $J$ is admissible for $K_{1},\ldots,K_{n}$, then for any $(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}$
$\lim\limits_{|t|\to\infty}\left(J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j})\right)=-\infty.$
Proof.
Let us prove the lemma for $t\to\infty$. If $t\to$-$\infty$, then we can consider $\widetilde{J}(t):=J(-t)$ and $\widetilde{K}_{j}(t):=K_{j}(-t),\ j=1,\ldots,n,$ and apply what is proved for $t\to\infty$.
1. Suppose that all $K^{\prime}_{j}$ be bounded below for large arguments. By our assumption and point 1 of Lemma 2.1, there are C,L>0C,\ L>0 such that
(11) $|K^{\prime}_{j}(t)|\leqslant C,\quad t\geqslant L,\ j=1,\ldots,n.$
Without loss of generality, assume that $\min\{t,t-y_{n}\}\geqslant L$. Using (7) and (11), we get
$K_{j}(t-y_{j})-K_{j}(t)\leqslant C|y_{j}|,\quad j=1,\ldots,n.$
Therefore, by the admissibility of $J$, we obtain
$J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j})\leqslant J(t)+\sum\limits_{j=1}^{n}(K_{j}(t)+C|y_{j}|)\to-\infty,\quad t\to\infty.$
2. Now assume that $K^{\prime}_{i}$ is not bounded below for some $i\in\{1,\ldots,n\}$ and for large $t.$ Denote
$f(\mathbf{y},t):=\sum\limits_{j=1}^{n}K_{j}(t-y_{j}).$
Note that $f(\mathbf{y},\cdot)$ is concave on $(y_{n},\infty)$. Take arbitrary point $t_{0}\in(y_{n},\infty)$ where $f^{\prime}_{t}$ exists. For any $t\in(y_{n},\infty)$, we have [17, p. 12, Th. D]
(12) $f(\mathbf{y},t)\leqslant f^{\prime}_{t}(\mathbf{y},t_{0})\cdot(t-t_{0})+f(\mathbf{y},t_{0}).$
By point 1 of Lemma 2.1 for $K^{\prime}_{j},\ j=1\ldots,n,$ and our assumption regarding $K_{i}$, we conclude that $f^{\prime}_{t}(\mathbf{y},t_{0})<0$ for large $t_{0}$ as a sum of non-increasing functions and a function not bounded below. Due to (12), we obtain
$\lim\limits_{t\to\infty}f(\mathbf{y},t)=-\infty.$
Since $J$ is bounded above, we finally get
$F(\mathbf{y},t)=J(t)+f(\mathbf{y},t)\to-\infty,\quad t\to\infty.$
3. Behavior of sums of translates for large arguments
In this section, we prove a key lemma which allows us to reduce our problem on the axis to the case of the segment. Another form of this statement was proven for sums of translates with positive multiples of a single kernel in [6, Lemma 4.3]. Now we are dealing with several kernels and we need a slightly modified estimate, so we will provide a proof.
The following statement is well-known, see, e.g., [6, Lemma 4.1].
Lemma A.
Suppose that a function $g$ is concave, nondecreasing on the semiaxis $M,\infty)$ and is continuous at $M.$ Then $g$ is uniformly continuous on $M,\infty)$.
We also need the following lemma.
Lemma 3.1.
If $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ are kernels and $J\colon\mathbb{R}\to\underline{\mathbb{R}}$ is an admissible field, then for any $\mathbf{y}\in\overline{S^{\infty}}$ we have
$\lim\limits_{\mathbf{x}\to\mathbf{y},\ |t|\to\infty}F(\mathbf{x},t)=-\infty.$
Proof.
Let us prove the lemma for $t\to\infty.$ For $t\to$-$\infty$ the proof is carried out by considering the reflection $\widetilde{K}(t):=K(-t),\ \widetilde{J}(t)=J(-t).$
Denote $I_{1}:=\{j:\ K_{j}\text{ is non-decreasing on }(0,\infty)\}.$ Let $j\in I_{1}.$ By Lemma A, the function $K_{j}$ is uniformly continuous on $1,\infty)$. Hence if $\|\mathbf{x}-\mathbf{y}\|$ is sufficiently small, then there exists $C_{j}>0$ such that for large $t$ we have $K_{j}(t-x_{j})-K_{j}(t-y_{j})\leqslant C_{j}$.
Now consider $I_{2}:=\{j:\ K_{j}\text{ is not non-decreasing on }(0,\infty)\}$ and $j\in I_{2}.$ By the concavity, $K_{j}$ decreases for large $t$. Since $\mathbf{x}$ converges, there is $c$ such that $x_{j}\leqslant c$ for all $j.$ So, for large $t$ we have that $K_{j}(t-x_{j})\leqslant K_{j}(t-c).$
Using the inequalities above, we get for large $t$ and for $\mathbf{x}$ sufficiently close to $\mathbf{y}$
$F(\mathbf{x},t)=J(t)+\sum\limits_{j=1}^{n}K_{j}(t-x_{j})\leqslant J(t)+\sum\limits_{j\in I_{1}}(K_{j}(t-y_{j})+C_{j})+\sum\limits_{j\in I_{2}}K_{j}(t-c).$
Applying Lemma 2.2, we obtain $\lim\limits_{\mathbf{x}\to\mathbf{y},\ t\to\infty}F(\mathbf{x},t)=-\infty.$
Now, let us prove the main lemma of this section.
Lemma 3.2.
Let $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be kernels and $J:\mathbb{R}\to\underline{\mathbb{R}}$ be an admissible field. Then for any $N\in\mathbb{N}$ there is a number $\tau_{N}>N$ such that for each $\mathbf{y}\in\overline{S^{N}}$
$\left(t<-\tau_{N}\implies F(\mathbf{y},t)\leqslant m_{0}(\mathbf{y})-1\right)\ \text{and}\ \left(t>\tau_{N}\implies F(\mathbf{y},t)\leqslant m_{n}(\mathbf{y})-1\right).$
In particular, this implies that
(13) $m^{\tau_{N}}_{j}(\mathbf{y})=m_{j}(\mathbf{y}),\quad\mathbf{y}\in\overline{S^{N}},\quad j=0,\ldots,n.$
Proof.
Let us prove the statement for $t<-\tau_{N}$. The proof of the second part is similar.
Assume for a contradiction that for some $N\in\mathbb{N}$
$\forall M\in\mathbb{N},\ M>N\quad\exists\,\mathbf{y}_{M}\in\overline{S^{N}}\quad\exists\,t_{M}\in\mathbb{R}\quad\left(t_{M}<-M\quad\&\quad F(\mathbf{y}_{M},t_{M})>m_{0}(\mathbf{y}_{M})-1\right).$
Take $x\in(-\infty,-N)$. By our assumption, we have for all $k$ with $M_{k}<x<-N$
$J(x)=F(\mathbf{y}_{M_{k}},x)-\sum\limits_{j=1}^{n}K_{j}(x-y^{M_{k}}_{j})\leqslant F(\mathbf{y}_{M_{k}},t_{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(x-y^{M_{k}}_{j}).$
Using continuity of $K$ at $x-y^{*}_{j}<0$ and Lemma 3.1, we get
$J(x)$ $\leqslant\lim\limits_{k\to\infty}\left(F(\mathbf{y}_{M_{k}},t_{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(x-y^{M_{k}}_{j})\right)$
$=\lim\limits_{k\to\infty}F(\mathbf{y}_{M_{k}},t_{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(x-y^{*}_{j})=-\infty.$
So, $J(x)\equiv$-$\infty$ for x<N.x<-N. We have a contradiction with our assumption, since then also $F(\mathbf{y}_{M},t_{M})=-\infty.$
4. Local homeomorphism
In this section, we prove a statement about local homeomorphism by reducing the problem to a segment. For $N\in\mathbb{N}$ take $\tau_{N}>N$ from Lemma 3.2. Consider
$K^{\tau_{N}}_{j}:=K_{j}|_{[-2\tau_{N},2\tau_{N}]},\quad J^{\tau_{N}}:=J|_{[-\tau_{N},\tau_{N}]}.$
Obviously, if $K_{j}$ is a singular (strictly) concave kernel function, then $K^{\tau_{N}}_{j}$ has the same properties. And it is also clear that if $J$ is an $n$-field function on $\mathbb{R}$, then $J^{\tau_{N}}$ is an $n$-field on $\tau_{N},\tau_{N}]$ for large $N$.
Note that by (13),
(14) $D^{\tau_{N}}(\mathbf{y})=D(\mathbf{y}),\quad\mathbf{y}\in R^{N},$
where $R^{N}$ is the regularity set for $F|_{[-N,N]}.$
Farkas, Nagy and Révész in [9] established the results mentioned in this section for sums of translates on the segment $[0,1]$. By applying the reasoning from Remark 1.3, these results can be extended to the segment $[a,b]$.
Lemma 4.1.
Let $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be singular kernels and $J:\mathbb{R}\to\underline{\mathbb{R}}$ be an admissible field. Then the regularity set $R$ is open and pathwise connected.
Proof.
Note that
$R=\bigcup\limits_{N\in\mathbb{N}}R^{N}.$
Farkas, Nagy and Révész proved in [9, Prop. 4.1] that the sets $R^{N}$ are open and pathwise connected. Hence $R$ is also open and pathwise connected as a union of sets with these properties. ∎
We will show that an analogue of the following property [9, Lemma 7.3] for $D^{[a,b]}$ holds for $D$.
Lemma B.
Let $K^{[a,b]}_{j}:(a-b,0)\cup(0,b-a)\to\underline{\mathbb{R}},\ j=1,\ldots,n,$ be singular kernels and $J^{[a,b]}:[a,b]\to\underline{\mathbb{R}}$ be an $n$-field function. Assume that for $j\in\{1,\ldots,n\}$
$\left(K^{[a,b]}_{j}(t)-K^{[a,b]}_{j}(t-(b-a))\right)^{\prime}\geqslant 0\quad\text{almost everywhere on $(0,b-a)$},$
and for any $\mathbf{y}\in R^{[a,b]}$ there exists $\eta>0$ such that
(15) $\text{either}\ F^{[a,b]}(\mathbf{y},t)$ $\leqslant m^{[a,b]}_{0}(\mathbf{y})-1,\ t\in[a,a+\eta]$
$\text{or}\ F^{[a,b]}(\mathbf{y},t)$ $\leqslant m^{[a,b]}_{n}(\mathbf{y})-1,\ t\in[b-\eta,b].$
Then the difference function $D^{[a,b]}$ is a local homeomorphism. Moreover, $D^{[a,b]}$ is locally bi-Lipschitz.
In fact, Farkas, Nagy and Révész assume conditions (3) instead of (15). In [9, Lemma 7.3] it is shown that (3) implies conditions (15), used in the proof of the local homeomorphism. For $\mathbb{R}$, conditions (15) in the problem on $\tau_{N},\tau_{N}]$ are consequences of Lemma 3.2. Note that the main role in the proof of Lemma 3.2 is played by admissibility of the field, similar to conditions (3).
Lemma 4.2.
Let $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be singular kernels satisfying ($GM$) and $J:\mathbb{R}\to\underline{\mathbb{R}}$ be an admissible field.
  1. (1)
    The difference function $D$ is a local homeomorphism. Moreover, $D$ is locally bi-Lipschitz.
  2. (2)
    The functions $m_{j}:\overline{S^{\infty}}\to\underline{\mathbb{R}},\ j=0,\ldots,n,$ are continuous in the extended sense.
Proof.
  1. (1)
    Let us apply Lemma B with $a,b]=[-\tau_{N},\tau_{N}]$ and
    $K^{\tau_{N}}_{j}:=K_{j}|_{[-2\tau_{N},2\tau_{N}]},\quad J^{\tau_{N}}:=J|_{[-\tau_{N},\tau_{N}]},\quad F^{\tau_{N}}:=F|_{[-\tau_{N},\tau_{N}]}$
    for large $N$. As discussed above, $K^{\tau_{N}}_{j}$ are singular kernels on $(-2\tau_{N},2\tau_{N})$, and $J^{\tau_{N}}$ is an $n$-field on $\tau_{N},\tau_{N}]$.
    Without loss of generality, we can assume that the statement of Lemma 3.2 is true for $\tau_{N}-1$. Then for $\mathbf{y}\in\overline{S^{N}}\cap R^{N}$ we get
    $F^{\tau_{N}}(\mathbf{y},t)\leqslant m^{\tau_{N}}_{0}(\mathbf{y})-1,\ t\in[-\tau_{N},-\tau_{N}+1]$
    and
    $F^{\tau_{N}}(\mathbf{y},t)\leqslant m^{\tau_{N}}_{n}(\mathbf{y})-1,\ t\in[\tau_{N}-1,\tau_{N}].$
    By (10) for $j=1,\ldots,n$ we have
    $\left(K^{\tau_{N}}_{j}(t)-K^{\tau_{N}}_{j}(t-2\tau_{N})\right)^{\prime}_{t}\geqslant 0\quad\text{almost everywhere on $(0,2\tau_{N})$}.$
    So, all conditions of Lemma B are satisfied. We obtain that the difference function $D^{\tau_{N}}$ is a local homeomorphism, and it is locally bi-Lipschitz on $R^{N}$. By (14), $D$ also has these properties on $R^{N}$ and thus on $R,$ since $N$ can be arbitrarily large.
  2. (2)
    In [9, Lemma 3.3] it is proven that the functions $m^{\tau_{N}}_{j},\ j=0,\ldots,n,$ are continuous on $\overline{S^{N}}$ in the extended sense. By (13), we conclude that $m_{j},\ j=0,\ldots,n,$ are extended continuous on $\overline{S^{N}}$, and hence on $\overline{S^{\infty}}$ as well due to the arbitrariness of NN.
5. Properness of the difference function
The idea of the proof of the following lemma was communicated to us by Szilárd Gy. Révész. We already used this idea in [6, Th. 3.1] with his permission. However, we now prove a statement of a different nature, and therefore we need a slightly different estimate in the proof. Moreover, we are dealing with several kernels satisfying ($GM$) instead of just one monotone kernel.
Lemma 5.1.
Assume that the kernels $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ satisfy (GMGM), and $J$ is an admissible field. Let $\{\mathbf{y}^{N}\}\subset R$ be an unbounded sequence convergent in the extended sense. Then there exists $i\in\{1,\ldots,n\}$ such that
$|m_{i}(\mathbf{y}^{N})-m_{i-1}(\mathbf{y}^{N})|\to\infty,\quad N\to\infty.$
Proof.
1. Denote $\mathbf{y}^{N}:=(y^{N}_{1},\ldots,y^{N}_{n}),\ y^{N}_{0}:=-\infty,\ y^{N}_{n+1}:=\infty.$ If $y^{N}_{n}\to\infty,$ take $i:=\min\{j:y^{N}_{j}\to\infty\}.$ Otherwise, we have $y^{N}_{1}\to$-$\infty$ and take $i:=\max\{j:y^{N}_{j}\to-\infty\}.$ Without loss of generality, consider the first case. The proof for the second case is similar.
2. Let us show that there exists $z>0$ such that $J(z)\neq$-$\infty$ and $z\in(y^{N}_{i-1},y^{N}_{i})$ for $N$ large enough.
By definition of $i$, we have either $y^{N}_{i-1}\to$-$\infty$ or $y^{N}_{i-1}\to y_{i-1}\in\mathbb{R}.$ In both cases, there exist $A>0,\ N_{0}\in\mathbb{N}$ such that $y^{N}_{i-1}<A-1$ for all $N>N_{0}.$ Since $\{\mathbf{y}^{N}\}\subset R$ for all $N$, there are arbitrarily large points where $J$ is finite. Fix one such point $z>A.$ Then, for all $N>N_{0}$, we have $y^{N}_{i-1}<z-1$. Moreover, there exists $N_{1}>N_{0}$ such that $y^{N}_{i}>z+1$ for all $N>N_{1}$. For $N>N_{1},$ we conclude that $z\in(y^{N}_{i-1},y^{N}_{i})$ and $|z-y^{N}_{j}|>1$ for all $j.$ Next we consider these $N$.
3. Let us estimate
$F(\mathbf{y}^{N},t)=J(t)+\sum\limits_{j<i}K_{j}(t-y^{N}_{j})+\sum\limits_{j\geqslant i}K_{j}(t-y^{N}_{j})$
for $t\in(y^{N}_{i},y^{N}_{i+1})$.
If $j<i,$ apply (8) with $t_{1}=1,\ t_{2}=t-z+1,\ h=z-1-y^{N}_{j}$. If $j>i$, use (9) with $t_{1}=z-y^{N}_{j},\ t_{2}=1,\ h=t-z.$ In both cases, we obtain
(16) $K_{j}(t-y^{N}_{j})-K_{j}(z-y^{N}_{j})\leqslant K_{j}(t-z+1)-K_{j}(1),\quad j\neq i.$
Since $K_{i}$ satisfies ($GM$), by (9) with $t_{1}=z-y^{N}_{i},\ t_{2}=t-y^{N}_{i},\ h=y^{N}_{i}-z-1$, we get
(17) $K_{i}(t-y^{N}_{i})-K_{i}(z-y^{N}_{i})\leqslant K_{i}(t-z-1)-K_{i}(-1).$
By (16) and (17), we obtain
$F(\mathbf{y}^{N},t)$ $\leqslant J(t)+\sum\limits_{j\neq i}(K_{j}(z-y^{N}_{j})+K_{j}(t-z+1)-K_{j}(1))$
$+K_{i}(z-y^{N}_{i})+K_{i}(t-z-1)-K_{i}(-1)$
$=F(\mathbf{y}^{N},z)+J(t)-J(z)+\sum\limits_{j\neq i}(K_{j}(t-z+1)-K_{j}(1))$
$+K_{i}(t-z-1)-K_{i}(-1).$
Let $C:=-J(z)-\sum\limits_{j\neq i}K_{j}(1)-K_{i}(-1).$ By the above estimate, we have
(18) $F(\mathbf{y}^{N},t)\leqslant F(\mathbf{y}^{N},z)+J(t)+\sum\limits_{j\neq i}K_{j}(t-z+1)+K_{i}(t-z-1)+C.$
Since $J$ is admissible, by Lemma 2.2, we have that
(19) $J(t)+\sum\limits_{j\neq i}K_{j}(t-z+1)+K_{i}(t-z-1)\to-\infty,\quad t\to\infty.$
Take arbitrary $L>0$. Since $t\geqslant y^{N}_{i}\to\infty,$ for large $N$ we obtain that (19) is less than LC-L-C. Hence, by (18), for large NN, we get that for $t\in(y^{N}_{i},y^{N}_{i+1})$
$F(\mathbf{y}^{N},t)\leqslant F(\mathbf{y}^{N},z)-L.$
Therefore, taking into account the choice of $z$, we have
$m_{i}(\mathbf{y}^{N})\leqslant F(\mathbf{y}^{N},z)-L\leqslant m_{i-1}(\mathbf{y}^{N})-L.$
Thus
$m_{i}(\mathbf{y}^{N})-m_{i-1}(\mathbf{y}^{N})\to-\infty,\quad N\to\infty.$
Definition 5.1.
A function $g:A\to B$ between two Hausdorff topological spaces is called proper if for any compact set $Q\subset B$ we have that $g^{-1}(Q)$ is also a compact set [19, p. 20].
Lemma 5.2.
Assume that the kernel functions $K_{j}\colon(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ are singular and satisfy (GMGM). Let a field function $J$ be admissible. Then the difference function $D:R\to\mathbb{R}^{n}$ is proper.
Proof.
Let $Q\subset\mathbb{R}^{n}$ be a compact set. We need to show that $D^{-1}(Q)$ is also a compact set, i. e., $D^{-1}(Q)$ is closed and bounded.
Let us prove that $D^{-1}(Q)$ is closed. By Lemma 4.2, $D$ is continuous. Hence, since $Q$ is closed, $D^{-1}(Q)$ is relatively closed in $R$, i. e.,
$D^{-1}(Q)=A\cap R,\quad\text{where $A$ is closed in $\mathbb{R}^{n}$}.$
Consider $\{\mathbf{y}^{N}\}\subset D^{-1}(Q)$ and let $\mathbf{y}^{N}\to\mathbf{y},\ N\to\infty.$ We have that $\mathbf{y}\in A,$ since $\{\mathbf{y}^{N}\}\subset A$ and $A$ is closed in $\mathbb{R}^{n}$. If we prove that $\mathbf{y}\in R$, it immediately follows that $D^{-1}(Q)$ is closed.
The field $J$ is finite at least at $n+1$ points, so there is $i\in\{0,\ldots,n\}$ such that $m_{i}(\mathbf{y})\in\mathbb{R}.$ Moreover, since $\{\mathbf{y}^{N}\}\subset D^{-1}(Q),$ there exists $C>0$ such that for all $N\in\mathbb{N}$
$|m_{j}(\mathbf{y}^{N})-m_{j-1}(\mathbf{y}^{N})|\leqslant C,\quad j=1,\ldots,n.$
By Corollary 4.2, the functions $m_{j}$ are continuous. Therefore, the differences $m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})$ also satisfy $|m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})|\leqslant C$ for all $j=1,\ldots,n$. Taking into account the finiteness of $m_{i}(\mathbf{y})$, this implies that $\mathbf{y}\in R.$
It remains to show that $D^{-1}(Q)$ is bounded. Suppose, contrary to our claim, that there exists an unbounded sequence $\{\mathbf{y}^{N}\}\subset D^{-1}(Q)$. Without loss of generality, let $\mathbf{y}^{N}$ converges in the extended sense. By Lemma 5.1, there is $i\in\{1,\ldots,n\}$ such that
$|m_{i}(\mathbf{y}^{N})-m_{i-1}(\mathbf{y}^{N})|\to\infty,\quad N\to\infty.$
Hence $Q\supset D(\{\mathbf{y}^{N}\})$ is not bounded. This contradicts the assumption that $Q$ is a compact set, and this finishes the proof. ∎
6. Proof of the main result
To prove Theorem 1.1, we need the following sufficient condition, due to C. W. Ho [20], that a local homeomorphism is a global one.
Theorem C.
Let $A,B$ be pathwise connected, Hausdorff topological spaces with $B$ simply connected. Let $f:A\to B$ be a proper local homeomorphism. Then $f$ is a global homeomorphism between $A$ and $B$.
Proof of Theorem 1.1.
By Lemma 4.1, the regular set $R$ is pathwise connected. The difference function
$D:R\to\mathbb{R}^{n},\ \mathbf{y}\mapsto(m_{1}(\mathbf{y})-m_{0}(\mathbf{y}),m_{2}(\mathbf{y})-m_{1}(\mathbf{y}),\ldots,m_{n}(\mathbf{y})-m_{n-1}(\mathbf{y}))$
is a local homeomorphism by Corollary 4.2 and it is proper by Lemma 5.2. Therefore, by Theorem C, we obtain that $D$ is a global homeomorphism. ∎
7. Homeomorphism theorem for sums of translates
on the semiaxis
Let $J^{+}$ be an $n$-field on $0,\infty)$, $K_{j}:(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be kernels. Admissibility of the field $J^{+}$ for $K_{1},\ldots,K_{n}$ can be defined as follows:
$\lim_{t\to\infty}\left(J^{+}(t)+\sum\limits_{j=1}^{n}K_{j}(t)\right)=-\infty.$
Denote
$\overline{S^{[0,\infty)}}:=\{(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}:\ 0\leqslant y_{1}\leqslant\ldots\leqslant y_{n}<\infty\}.$
Consider sums of translates on $0,\infty)$
$F^{+}(\mathbf{y},t):=J^{+}(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j}),\quad\mathbf{y}\in\overline{S^{[0,\infty)}},\quad t\in[0,\infty),$
and local maxima
$m^{+}_{0}(\mathbf{y})$ $:=\sup\limits_{t\in[0,y_{1}]}F^{+}(\mathbf{y},t),\quad m^{+}_{n}(\mathbf{y}):=\sup\limits_{t\in[y_{n},\infty)}F^{+}(\mathbf{y},t),$
$m^{+}_{j}(\mathbf{y})$ $:=\sup\limits_{t\in[y_{j},y_{j+1}]}F^{+}(\mathbf{y},t),\quad j=1,\ldots,n-1.$
Let
$R^{+}:=\{\mathbf{y}\in\overline{S^{[0,\infty)}}:\ m^{+}_{j}(\mathbf{y})\neq-\infty\text{ for }j=0,\ldots,n\}.$
Corollary 7.1.
Suppose that the singular, strictly concave kernel functions $K_{j}:(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ satisfy ($GM$) and $J^{+}:[0,\infty)\to\underline{\mathbb{R}}$ is an admissible $n$-field function for $K_{1},\ldots,K_{n}$. Then the difference function
$D^{+}:R^{+}\to\mathbb{R}^{n},\quad\mathbf{y}\mapsto(m^{+}_{1}(\mathbf{y})-m^{+}_{0}(\mathbf{y}),m^{+}_{2}(\mathbf{y})-m^{+}_{1}(\mathbf{y}),\ldots,m^{+}_{n}(\mathbf{y})-m^{+}_{n-1}(\mathbf{y}))$
is a homeomorphism between $R^{+}$ and $\mathbb{R}^{n}$. Moreover, $D^{+}$ is locally bi-Lipschitz.
Proof.
Consider
$F(\mathbf{y},t)=J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j}),\quad\mathbf{y}\in\overline{S^{[0,\infty)}},\quad t\in\mathbb{R},$
where
$J(t)=\begin{cases}-\infty,&t<0,\\ J^{+}(t),&t\geqslant 0.\end{cases}$
Note that $K_{j},\ j=1,\ldots,n,$ and $J$ satisfy the conditions of Theorem 1.1. Therefore, the difference function $D:R\to\mathbb{R}^{n}$ for $F$ is a homeomorphism, and it is locally bi-Lipschitz. It remains to note that
$R^{+}=R,\quad\text{and}\quad m_{j}^{+}(\mathbf{y})=m_{j}(\mathbf{y}),\ \mathbf{y}\in R^{+}.$
Hence $D^{+}\equiv D,$ and $D^{+}$ is a homeomorphism, and it is locally bi-Lipschitz, too. ∎
8. Homeomorphism theorem for weighted generalized
polynomials on the real axis
Let $r_{1},\ldots,r_{n}>0$ be arbitrary. Denote $\mathbf{r}:=(r_{1},\ldots,r_{n})$. Consider the following set of monic generalized nonnegative polynomials [21, p. 392] of degree $r\leavevmode\nobreak\ :=\leavevmode\nobreak\ \sum\limits_{j=1}^{n}r_{j}$
$\mathcal{P}_{\mathbf{r}}(\mathbb{R}):=\left\{p(\mathbf{y},t)=\prod\limits_{j=1}^{n}|t-y_{j}|^{r_{j}}:\ -\infty<y_{1}<\ldots<y_{n}<\infty\right\}.$
Let $\mathbb{R}^{n}_{+}$ denote the subset of vectors from $\mathbb{R}^{n}$ with positive coordinates.
Corollary 8.1.
Let $w:\mathbb{R}\to[0,\infty)$ be a bounded above function assuming non-zero values at more than $n$ points. Assume that
$\lim\limits_{|t|\to\infty}w(t)\cdot t^{r}=0.$
Let
$S^{\infty}:=\{(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}:\ -\infty<y_{1}<\ldots<y_{n}<\infty\},\quad$
and
$X:=\{t\in\mathbb{R}:\ w(t)=0\}.$
For convenience, let $y_{0}:=-\infty,\ y_{n+1}:=\infty.$ Denote
$R:=\{\mathbf{y}\in S^{\infty}:\ (y_{j},y_{j+1})\not\subseteq X\quad\text{for}\quad j=0,\ldots,n\}.$
Then the mapping
$R\ni\mathbf{y}\mapsto\left(\dfrac{\sup\limits_{t\in(y_{1},y_{2})}w(t)p(\mathbf{y},t)}{\sup\limits_{t\in(-\infty,y_{1})}w(t)p(\mathbf{y},t)},\ldots,\dfrac{\sup\limits_{t\in(y_{n},\infty)}w(t)p(\mathbf{y},t)}{\sup\limits_{t\in(y_{n-1},y_{n})}w(t)p(\mathbf{y},t)}\right)\in\mathbb{R}^{n}_{+}$
is a homeomorphism between $R$ and $\mathbb{R}^{n}_{+}$. Moreover, it is locally bi-Lipschitz.
Proof.
Let $K_{j}:=r_{j}\log|\cdot|,\ j=1,\ldots,n,\ J:=\log w.$ Obviously, these kernels and field satisfy the conditions of Theorem 1.1. Therefore, the difference function
$D:R\to\mathbb{R}^{n},\quad\mathbf{y}\mapsto\left(\sup\limits_{t\in(y_{j},y_{j+1})}\log(w(t)p(\mathbf{y},t))-\sup\limits_{t\in(y_{j-1},y_{j})}\log(w(t)p(\mathbf{y},t))\right)_{j=1}^{n}$
is a homeomorphism between $R$ and $\mathbb{R}^{n}$, and it is locally bi-Lipschitz. Since the logarithm strictly increases on $(0,\infty),$
$\sup\limits_{t\in(y_{j},y_{j+1})}\log(w(t)p(\mathbf{y},t))=\log\left(\sup\limits_{t\in(y_{j},y_{j+1})}w(t)p(\mathbf{y},t)\right),\quad\mathbf{y}\in R.$
Let us write the difference function $D$ as
$\mathbf{y}\mapsto\left(\log\dfrac{\sup\limits_{t\in(y_{j},y_{j+1})}w(t)p(\mathbf{y},t)}{\sup\limits_{t\in(y_{j-1},y_{j})}w(t)p(\mathbf{y},t)}\right)_{j=1}^{n}$
Note that the function
[…]
is a homeomorphism, and it is locally Lipschitz. It remains to see that the mapping in our statement is equal to $E\circ D.$
9. Interpolation by products of log-concave functions
with weight
The homeomorphism theorem allows one to obtain results on interpolation by weighted products of log-concave functions. This is discussed in detail in [9, Sect. 9] for the homeomorphism theorem on the segment. The authors proved general results [9, Th. 9.2, 9.6] on the existence and uniqueness of an interpolation function. They also studied applications of these results to trigonometric interpolation [9, Th. 9.7]. Moreover, in [9, Subsect. 9.3, 9.4] the authors investigated the applications to moving node Hermite–Fejér interpolation. Below we prove some analogues of the most general of these results on the real axis.
9.1. Abstract log-concave interpolation on the real axis
Theorem 9.1.
Let $L_{1},\ldots,L_{n}:\mathbb{R}\to[0,\infty)$ be log-concave functions vanishing at 0 and satisfying
$\lim\limits_{t\to-\infty}(\log L_{j}(t))^{\prime}\leqslant\lim\limits_{t\to\infty}(\log L_{j}(t))^{\prime},\quad j=1,\ldots,n.$
Let $w:\mathbb{R}\to[0,\infty)$ be a bounded above function assuming non-zero values at more than $n$ points. Assume that
$\lim\limits_{|t|\to\infty}w(t)\prod\limits_{j=1}^{n}L_{j}(t)=0.$
For any $\infty<x_{0}<\ldots<x_{n}<\infty$ with $w(x_{j})>0$ and $\alpha_{0},\ldots,\alpha_{n}>0$ there are a unique $C>0$ and points $y_{1}<y_{2}<\ldots<y_{n}$ with $x_{j}<y_{j+1}<x_{j+1}$ for each $j\in\{0,\ldots,n-1\}$ such that for the function
$G(t):=Cw(t)\prod\limits_{j=1}^{n}L_{j}(t-y_{j})$
we have
$G(x_{j})=\alpha_{j},\quad j=0,\ldots,n.$
Proof.
Let
$K_{j}(t)$ $:=\log L_{j}(t),\quad j=1,\ldots,n,$
$J(t)$ $:=\begin{cases}\log w(x_{j}),\quad t=x_{j},\ j=0,\ldots,n,\\ -\infty,\quad t\in\mathbb{R}\setminus\{x_{0},\ldots,x_{n}\}.\end{cases}$
It is easy to see that $K_{j}$ and $J$ satisfy the conditions of Theorem 1.1.
By construction of $J$, the regularity set has the following form
$R:=\{\mathbf{y}\in\overline{S^{\infty}}:\ x_{j}<y_{j+1}<x_{j+1}\text{ for }j=0,\ldots,n-1\}.$
Take arbitrary $\alpha_{0},\ldots,\alpha_{n}>0.$ By Theorem 1.1, there is a unique $\mathbf{y}\in R$ such that
(20) $m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})=\log(\alpha_{j}/\alpha_{j-1}),\quad j=1,\ldots,n.$
We have
(21) $m_{j}(\mathbf{y})=F(\mathbf{y},x_{j})=\log w(x_{j})+\sum\limits_{k=1}^{n}\log L_{k}(x_{j}-y_{k}),\quad j=0,\ldots,n.$
Therefore, using (20), after exponentiating we get
$\dfrac{\exp m_{j}(\mathbf{y})}{\exp m_{j-1}(\mathbf{y})}=\dfrac{w(x_{j})\prod\limits_{k=1}^{n}L_{k}(x_{j}-y_{k})}{w(x_{j-1})\prod\limits_{k=1}^{n}L_{k}(x_{j-1}-y_{k})}=\dfrac{\alpha_{j}}{\alpha_{j-1}},\quad j=1,\ldots,n.$
Hence
$\dfrac{w(x_{j})\prod\limits_{k=1}^{n}L_{k}(x_{j}-y_{k})}{w(x_{0})\prod\limits_{k=1}^{n}L_{k}(x_{0}-y_{k})}=\dfrac{\alpha_{j}}{\alpha_{0}},\quad j=1,\ldots,n,$
and for the function
$G(t):=Cw(t)\prod\limits_{k=1}^{n}L_{k}(t-y_{k})\quad\text{with}\quad C:=\dfrac{\alpha_{0}}{w(x_{0})\prod\limits_{k=1}^{n}L_{k}(x_{0}-y_{k})},$
we obtain
$G(x_{j})=\alpha_{j}\quad j=0,\ldots,n.$
Let us show the uniqueness. Assume that for some $\widetilde{C}>0$ and $\widetilde{\mathbf{y}}=(\widetilde{y}_{1},\ldots,\widetilde{y}_{n})\in R$ we have
$\widetilde{C}w(x_{j})\prod\limits_{k=1}^{n}L_{k}(x_{j}-\widetilde{y}_{k})=\alpha_{j}\quad j=0,\ldots,n.$
Using (21), we have
$\log\widetilde{C}+m_{j}(\widetilde{\mathbf{y}})=\log\widetilde{C}+\log w(x_{j})+\sum\limits_{k=1}^{n}\log L_{k}(x_{j}-y_{k})=\log\alpha_{j}\quad j=0,\ldots,n.$
Hence
$m_{j}(\widetilde{\mathbf{y}})-m_{j-1}(\widetilde{\mathbf{y}})=\log(\alpha_{j}/\alpha_{j-1}),\quad j=1,\ldots,n.$
By Theorem 1.1, we conclude that $\widetilde{\mathbf{y}}=\mathbf{y},$ therefore, $\widetilde{C}=C,$ too. ∎
9.2. Moving node Hermite–Fejér interpolation
Theorem 9.2.
Let $L_{1},\ldots,L_{n}:\mathbb{R}\to[0,\infty)$ be strictly log-concave functions vanishing at 0 and satisfying
$\lim\limits_{t\to-\infty}(\log L_{j}(t))^{\prime}\leqslant\lim\limits_{t\to\infty}(\log L_{j}(t))^{\prime},\quad j=1,\ldots,n.$
Let $w:\mathbb{R}\to[0,\infty)$ be an upper semicontinuous function assuming non-zero values at more than $n$ points. Assume that
$\lim\limits_{|t|\to\infty}w(t)\prod\limits_{j=1}^{n}L_{j}(t)=0.$
For convenience, let $y_{0}:=$-$\infty$ and $y_{n+1}:=\infty$. For any $\alpha_{0},\ldots,\alpha_{n}>0$ there are a unique $C>0$ and points $y_{1}<\ldots<y_{n}$ such that for the function
$G(t):=Cw(t)\prod\limits_{j=1}^{n}L_{j}(t-y_{j})$
there are $z_{0},\ldots,z_{n}$ with $z_{0}<y_{1}<z_{1}<y_{2}<\ldots<z_{n-1}<y_{n}<z_{n}$ and
$G(z_{j})=\alpha_{j},\quad j=0,\ldots,n,$
where $z_{j}$ is the maximum point of $G$ between $y_{j}$ and $y_{j+1}$ for each $j=0,\ldots,n.$
Proof.
Let
$K_{j}(t):=\log L_{j}(t),\quad j=1,\ldots,n,\quad J(t):=\log w(t).$
By Theorem 1.1, there is a unique $\mathbf{y}\in R$ such that
$m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})=\log(\alpha_{j}/\alpha_{j-1}),\quad j=1,\ldots,n.$
Note that since $K_{j}$ are strictly concave and $w$ is upper semicontinuous, there are $z_{0},\ldots,z_{n}$ with $z_{0}<y_{1}<z_{1}<y_{2}<\ldots<z_{n-1}<y_{n}<z_{n}$ such that
$F(\mathbf{y},z_{i})=m_{i}(\mathbf{y}),\quad i=0,\ldots,n.$
The further proof is similar to the proof of Theorem 9.1 with $x_{j}$ replaced by $z_{j},\\ j=1,\ldots,n$. ∎
Acknowledgements
The author is grateful to Prof. Sz. Gy. Révész for proposing the problem, useful references, discussions and detailed feedback during the preparation of the article.
The author is thankful to P. Yu. Glazyrina for continuous support, discussions on the proofs, and helpful suggestions regarding the structure of the article.